OpenStudy (anonymous):

7sqrt 4x+2sqrt25x+3sqrt16=

OpenStudy (anonymous):

\[7\sqrt {4x}+2\sqrt{25x}+3\sqrt{16}\] \[=7\sqrt{2*2*x}+\sqrt{5*5*x}+3\sqrt{4*4}\] \[=14\sqrt{x}+10\sqrt{x}+12\] \[=24\sqrt{x}+12\] \[=12(\sqrt{x}+1)\]

OpenStudy (anonymous):

sorry missed a 2 in the ans answer is \[=12(2\sqrt{x}+1)\]